# On-Device Multimodal Federated Learning for Efficient Jamming Detection

### v1

@@ Abstract
Wireless networks face severe vulnerabilities from jamming attacks, which can significantly disrupt communication. Existing detection approaches are often unimodal, rely on centralized processing, and demand substantial computational resources, hindering scalability, efficiency, and deployment feasibility. To address these challenges, we introduce a multimodal Federated Learning (FL) framework for on-device jamming detection and classification that integrates spectrograms with cross-layer network Key Performance Indicators (KPIs) through a lightweight dual-encoder architecture equipped with a fusion module and a multimodal projection head. This design enables privacy-preserving training and inference by ensuring that only model parameters are exchanged, while raw data remains on the device. The framework is implemented and evaluated on a wireless experimental testbed using, to the best of our knowledge, the first over-the-air multimodal dataset with synchronized benign and three distinct jamming scenarios. Results show that our approach surpasses state-of-the-art unimodal baselines by up to 15% in detection accuracy, achieves convergence with 60% fewer communication rounds, and maintains low resource usage. Its benefits are most evident under heterogeneous data distributions across devices, where it exhibits strong robustness and reliability.
Index Terms:
Security, Jamming Detection, Multimodal Federated Learning, Edge Intelligence, On-Device Inference.
@@ I Introduction
Wireless networks serve as the backbone of modern communication systems, enabling a wide range of services. However, their open and shared medium exposes them to physical-layer security threats; particularly jamming attacks, where adversaries intentionally interfere with legitimate wireless signals to disrupt communication [1]. Once largely associated with military operations, jamming has emerged as a broader threat vector, with recent incidents disrupting commercial aviation via GPS interference [2] and disabling home security systems through Wi-Fi jamming during burglaries [3]. The accessibility of low-cost Software-Defined Radios (SDRs) and open-source software has drastically lowered the barrier to launching such attacks. For instance, USB dongles costing under \$20 can be configured to emit disruptive signals, enabling denial-of-service attacks with minimal technical expertise [4].
In response, several techniques have been introduced to detect and categorize such attack vectors. Conventional approaches rely on energy-based methods, monitoring signal strength anomalies or disruptions in Key Performance Indicators (KPIs) [1], while state-of-the-art approaches increasingly rely on Machine Learning (ML) to analyze In-phase and Quadrature (I/Q) samples and detect interference patterns. For instance, common approaches transform I/Q streams into spectrogram representations and employ computer vision models to detect jamming activity  [5, 6].
Despite their demonstrated effectiveness, existing ML-based methods continue to face challenges in terms of computational efficiency, scalability, and real-world deployability. First, many proposed detection models are computationally intensive, employing deep and complex architectures that prioritize accuracy over efficiency, thereby constraining their deployability in resource-constrained environments [7, 8]. Second, most prior systems are designed as centralized solutions [9, 10, 11] that rely on collecting large volumes of raw data (e.g., I/Q streams) from sensors or base stations and transmitting them to a central server for analysis and decision-making. This process not only imposes substantial burdens on compute and network resources, but also raises scalability concerns. In addition, transmitting such data to centralized servers introduces additional security risks, as raw I/Q samples and KPIs may reveal sensitive information, including user identities, locations, and behavioral patterns [12]. Centralization further amplifies the risk of interception, tampering, and adversarial attacks, such as model poisoning [13].
@@
Beyond these challenges, prior detection solutions focus solely on unimodal sources, overlooking complementary information naturally present within the network across diverse modalities, which can limit their ability to form a holistic view of the network state [9, 11, 14]. This lack of cross-modal integration can result to misdetections, especially when sophisticated attackers mask their activity in one modality while leaving subtle traces in another [15]. Although multimodal architectures have demonstrated state-of-the-art performance in multiple domains such as vision-language understanding [16], their application to wireless security remains largely unexplored. A major obstacle is the lack of publicly available datasets that capture time-aligned multimodal network information, which are essential for developing and evaluating effective multimodal detection systems. Existing datasets are predominantly synthetic and fail to reflect real-world complexity, limiting the robustness, practical relevance, and deployment feasibility of current approaches [7, 8].
@@ II Related Work
Centralized Detection Approaches. Centralized ML-based detection systems have dominated early research on jamming detection and classification [9, 10, 5, 6, 7, 8, 17, 18, 19]. These approaches typically rely on offloading continuous, high-volume data, such as raw I/Q samples or KPI streams, to a centralized server for training and inference [17, 18, 11, 14]. To improve detection accuracy, many of these systems employ deep and complex models, such as computer vision-based architectures [9, 10, 5, 6, 7, 8]. Despite their high accuracy, these centralized solutions remain challenging to deploy in practice due to their significant compute and communication demands, as well as data privacy concerns.
@@
Distributed and Collaborative Detection Approaches. To improve scalability and preserve data privacy, recent efforts have explored FL-based techniques [20, 21, 22, 23, 24, 25]. These approaches enable multiple devices to collaboratively train and fine-tune a shared detection model, while also supporting local inference without exposing raw private data to an untrusted centralized entity. However, they still suffer from key limitations. Existing approaches rely on heavyweight ML models for jamming detection, assuming abundant computational resources—often unrealistic in IoT and mobile scenarios. Moreover, they depend on single-source inputs, overlooking complementary data such as time-series metrics that offer critical context for accurate detection.
@@
Unimodal Detection Approaches. Ongoing research on ML-based jamming detection has predominantly relied on unimodal learning, using either spectrum-level representations [9, 10, 5, 6, 7, 8, 17, 18, 19] or network-level statistics [11, 14, 26, 27]. Spectrum-centric approaches transform raw I/Q samples into spectrograms and employ deep learning models [9, 10, 5, 6, 7, 8] to identify interference patterns. In contrast, models based on network- or device-level statistics, such as signal quality, block error rate, and packet loss ratio, typically rely on threshold-based logic or anomaly detection frameworks that learn the normal behavior of the network and flag deviations [11, 14, 26, 27]. However, relying on a unimodal perspective limits these approaches’ ability to capture a comprehensive view of the network, reducing their effectiveness against sophisticated attacks.
@@
Multimodal Federated Learning. Multimodal FL integrates heterogeneous data sources, such as images, audio, and time-series signals, to capture complementary information and improve collaborative learning robustness. While this paradigm is gaining traction in fields such as healthcare, human activity recognition, and autonomous systems, its application to wireless security remains limited [28]. Despite its potential, existing multimodal FL approaches across domains still face two significant limitations. Firstly, they exhibit poor performance under non-IID conditions, an inherent challenge in real-world deployments where clients experience diverse environments and modality distributions [29]. Secondly, they frequently overlook the increased computational cost of processing multiple data streams, often assuming that all clients have uniform and sufficient resources [30]. These limitations hinder their applicability in resource-constrained, heterogeneous edge scenarios, such as wireless jamming detection.
@@ III-C Training Workflow
FedJam adopts a synchronous FL setup to collaboratively train a multimodal model across multiple edge clients. Each client performs local updates on its private spectrogram and KPI data, and periodically transmits selected model parameters, namely those of the spectrum encoder, KPI encoder, and classifier, to a central server. To minimize communication costs, these updates may be compressed before transmission. The server aggregates the received updates using any aggregation method (e.g., FedAvg [31]) and broadcasts the updated global model back to all clients. This process is repeated for a fixed number of communication rounds or until convergence, enabling the multimodal model to learn effectively from decentralized data without sharing raw samples.
@@ IV-A Experimental Environment
Network Connectivity. Connectivity between the edge devices and the central server is supported through both Wi-Fi and 5G. Wi-Fi connectivity is provided by a TP-Link Archer C7 router running OpenWRT firmware, operating exclusively in the 2.4 GHz band across the three non-overlapping 20 MHz channels (1, 6, and 11). 5G connectivity is enabled through a fully customized private 5G network operating in the n78 TDD band with 20 MHz bandwidth centered at 3.319 GHz, using a 5 ms radio frame structure with 7 downlink slots and 2 uplink slots at 30 kHz subcarrier spacing. This private 5G deployment is supported by a separate testbed developed in-house, shown in Fig. 3(b) and Fig. 3(c), featuring a monolithic base station (gNB), connected to a USRP N300 SDR via a 10 Gbps fronthaul link, with the SDR further connected to 6 dBi directional antennas to enable over-the-air transmission. The gNB is further connected to the 5G core network via a dedicated 10 Gbps backhaul link, with all 3GPP-compliant Virtual Network Functions (VNFs) deployed as containerized services within a customized Kubernetes cluster. Both the Radio Access Network (RAN) and core network VNFs are implemented using the latest release of the OpenAirInterface (OAI) 5G software stack [32].
@@ V-A Baselines
  • FewShotDense: A unimodal spectrogram-based convolutional neural network built on DenseNet121, previously used for few-shot interference classification tasks [25].
  • SpectroNet: A unimodal EfficientNet-B0-based architecture trained on spectrogram data for multi-class jamming detection and classification [9].
  • JamShield: A unimodal LSTM-based architecture trained on cross-layer KPI time series data for multi-class jamming detection [33].
@@
  • Notable Unimodal Models: We also experiment with two lightweight unimodal architectures to assess individual modality performance under constrained resources: (i) an 1D-CNN [34] operating on KPI time series and (ii) TinyViT [35], a compact vision transformer applied to spectrogram inputs.
@@ V-B FedJam Configuration
For spectral feature extraction, the Spectrum Encoder employs MobileNetV3-Small [36], a lightweight convolutional neural network pretrained on ImageNet, producing a 1024-dimensional feature vector that captures localized interference patterns in the frequency domain. The KPI Encoder is implemented using a compact Transformer-based architecture composed of four self-attention layers, preceded by a linear projection to 128 dimensions. The resulting sequence is aggregated using an attention-based pooling mechanism to produce a 128-dimensional embedding summarizing the network’s temporal performance. The Fusion Module combines the two modality-specific embeddings via concatenation, forming a 1152-dimensional joint representation. This fused vector is then passed to the Multimodal Projection Head Module, implemented as a two-layer fully connected neural network with 256 hidden units and ReLU activation, followed by an output layer producing logits over the four jamming classes. Finally, all of FedJam’s components are jointly fine-tuned end-to-end using our collected multimodal dataset.
@@ V-D Evaluation Results
For our simulation experiments, we utilized the Flower framework [37], a widely adopted and flexible platform in the FL community that enables evaluation of federated solutions across diverse scales and heterogeneous settings.
@@
FewShotDense [25] 95.10 $\pm$ 0.31 88.10 $\pm$ 0.24 84.10 $\pm$ 0.39
SpectroNet [9] 96.00 $\pm$ 0.27 93.80 $\pm$ 0.34 89.00 $\pm$ 0.25
JamShield [33] 58.40 $\pm$ 0.40 50.70 $\pm$ 0.21 46.70 $\pm$ 0.28
1D-CNN [34] 46.60 $\pm$ 0.33 42.53 $\pm$ 0.30 40.89 $\pm$ 0.29
TinyViT [35] 89.60 $\pm$ 0.87 84.60 $\pm$ 0.52 77.10 $\pm$ 0.92
@@ 10 Clients 25 Clients 50 Clients
In addition to fusion strategies, the choice of aggregation algorithm can influence convergence and robustness in FL. We evaluate three widely used aggregation methods: FedAvg, FedOpt [38], and FedProx [39], to assess their compatibility with our framework. As shown in Fig. 7(b), all algorithms perform similarly, with FedAvg offering a marginal improvement. These results highlight that our method does not rely on a specific aggregation scheme. Indeed, FedJam is designed to be modular and aggregation-agnostic, enabling seamless integration with future or domain-specific optimizers, and supporting flexible experimentation in practical deployments.
@@
We additionally evaluate the total training time required under different device energy modes, where all participating clients operate at the same power level. The Jetson Orin Nano devices support three energy modes: Mode 0 (15W), Mode 1 (25W), and Mode 2 (MAXN SUPER), which progressively scale CPU/GPU frequencies and memory bandwidth [40]. As shown in Fig. 11, higher energy modes consistently reduce training time across both Wi-Fi and 5G. For instance, under Wi-Fi with 4 clients, total training time decreases from 575 minutes (15W) to 451 minutes (MAXN SUPER), corresponding to a 21.6% reduction. Similarly, with 16 clients, training time drops from 290 to 220 minutes—a 24.1% improvement. These improvements reflect lower per-device training time due to increased hardware utilization and faster computation, although diminishing gains are observed beyond 8 clients due to reduced per-client workload.

### v2

@@ Abstract
Jamming attacks pose a critical threat to wireless networks, yet existing detection methods remain largely unimodal, centralized and resource-intensive, limiting their performance, scalability, and deployment feasibility, respectively. To address these limitations, we present FedJam, a multimodal Federated Learning (FL) framework for on-device jamming detection and classification. FedJam locally fuses spectrograms and cross-layer network Key Performance Indicators (KPIs) using a lightweight dual-encoder architecture with an integrated fusion module and multimodal projection head, that enables privacy-preserving training and inference without transmitting raw data. We prototype and deploy FedJam on a wireless experimental testbed and evaluate it using the first, over-the-air multimodal dataset comprising synchronized samples across benign and three distinct jamming attack types. FedJam outperforms state-of-the-art unimodal baselines by up to 15% in accuracy, while requiring 60% fewer communication rounds to converge, and maintains low resource utilization. Its advantage is especially pronounced in realistic scenarios, where it remains extremely robust under heterogeneous data distributions across devices.
@@ I Introduction
Wireless networks serve as the backbone of modern communication systems, enabling a wide range of services. However, their open and shared medium exposes them to physical-layer security threats; particularly jamming attacks, where adversaries intentionally interfere with legitimate wireless signals to disrupt communication [jamming_survey]. Once largely associated with military operations, jamming has emerged as a broader threat vector, with recent incidents disrupting commercial aviation via GPS interference [gps_jamming_baltic] and disabling home security systems through Wi-Fi jamming during burglaries [wifi_jamming_glendale]. The accessibility of low-cost Software-Defined Radios (SDRs) and open-source software has drastically lowered the barrier to launching such attacks. For instance, USB dongles costing under \$20 can be configured to emit disruptive signals, enabling denial-of-service attacks with minimal technical expertise [cheap_sdr_jamming].
In response, several techniques have been introduced to detect and categorize such attack vectors. Conventional approaches rely on energy-based methods, monitoring signal strength anomalies or disruptions in Key Performance Indicators (KPIs) [jamming_survey], while state-of-the-art approaches increasingly rely on Machine Learning (ML) to analyze In-phase and Quadrature (I/Q) samples and detect interference patterns. For instance, common approaches transform I/Q streams into spectrogram representations and employ computer vision models to detect jamming activity  [spectrogram_3, spectrogram_4].
Despite their demonstrated effectiveness, existing ML-based methods continue to face challenges in terms of computational efficiency, scalability, and real-world deployability. First, many proposed detection models are computationally intensive, employing deep and complex architectures that prioritize accuracy over efficiency, thereby constraining their deployability in resource-constrained environments [spectrogram_5, spectrogram_6]. Second, most prior systems are designed as centralized solutions [spectrogram_1, spectrogram_2, kpi1] that rely on collecting large volumes of raw data (e.g., I/Q streams) from sensors or base stations and transmitting them to a central server for analysis and decision-making. This process not only imposes substantial burdens on compute and network resources, but also raises scalability concerns. In addition, transmitting such data to centralized servers introduces additional security risks, as raw I/Q samples and KPIs may reveal sensitive information, including user identities, locations, and behavioral patterns [iq_security]. Centralization further amplifies the risk of interception, tampering, and adversarial attacks, such as model poisoning [adversarial_attacks].
@@
Beyond these challenges, prior detection solutions focus solely on unimodal sources, overlooking complementary information naturally present within the network across diverse modalities, which can limit their ability to form a holistic view of the network state [spectrogram_1, kpi1, kpi2]. This lack of cross-modal integration can result to misdetections, especially when sophisticated attackers mask their activity in one modality while leaving subtle traces in another [sumalatha2024comprehensive]. Although multimodal architectures have demonstrated state-of-the-art performance in multiple domains such as vision-language understanding [multimodal_sota], their application to wireless security remains largely unexplored. A major obstacle is the lack of publicly available datasets that capture time-aligned multimodal network information, which are essential for developing and evaluating effective multimodal detection systems. Existing datasets are predominantly synthetic and fail to reflect real-world complexity, limiting the robustness, practical relevance, and deployment feasibility of current approaches [spectrogram_5, spectrogram_6].
@@
  • We release all artifacts (including dataset, code, and deployment scripts)1 to foster reproducibility and enable further research.
@@ II Related Work
Centralized Detection Approaches. Centralized ML-based detection systems have dominated early research on jamming detection and classification [spectrogram_1, spectrogram_2, spectrogram_3, spectrogram_4, spectrogram_5, spectrogram_6, iq_samples_1_deep_learning, iq_samples_2_deep_learning, iq_samples_3_deep_learning]. These approaches typically rely on offloading continuous, high-volume data, such as raw I/Q samples or KPI streams, to a centralized server for training and inference [iq_samples_1_deep_learning, iq_samples_2_deep_learning, kpi1, kpi2]. To improve detection accuracy, many of these systems employ deep and complex models, such as computer vision-based architectures [spectrogram_1, spectrogram_2, spectrogram_3, spectrogram_4, spectrogram_5, spectrogram_6]. Despite their high accuracy, these centralized solutions remain challenging to deploy in practice due to their significant compute and communication demands, as well as data privacy concerns.
@@
Distributed and Collaborative Detection Approaches. To improve scalability and preserve data privacy, recent efforts have explored FL-based techniques [federated_learning_attacks, federated_learning_spectrogram_1, federated_learning_spectrogram_2, federated_learning_spectrogram_3, federated_learning_spectrogram_4, federated_learning_spectrogram_5]. These approaches enable multiple devices to collaboratively train and fine-tune a shared detection model, while also supporting local inference without exposing raw private data to an untrusted centralized entity. However, they still suffer from key limitations. Existing approaches rely on heavyweight ML models for jamming detection, assuming abundant computational resources—often unrealistic in IoT and mobile scenarios. Moreover, they depend on single-source inputs, overlooking complementary data such as time-series metrics that offer critical context for accurate detection.
@@
Unimodal Detection Approaches. Ongoing research on ML-based jamming detection has predominantly relied on unimodal learning, using either spectrum-level representations [spectrogram_1, spectrogram_2, spectrogram_3, spectrogram_4, spectrogram_5, spectrogram_6, iq_samples_1_deep_learning, iq_samples_2_deep_learning, iq_samples_3_deep_learning] or network-level statistics [kpi1, kpi2, kpi3, kpi4]. Spectrum-centric approaches transform raw I/Q samples into spectrograms and employ deep learning models [spectrogram_1, spectrogram_2, spectrogram_3, spectrogram_4, spectrogram_5, spectrogram_6] to identify interference patterns. In contrast, models based on network- or device-level statistics, such as signal quality, block error rate, and packet loss ratio, typically rely on threshold-based logic or anomaly detection frameworks that learn the normal behavior of the network and flag deviations [kpi1, kpi2, kpi3, kpi4]. However, relying on a unimodal perspective limits these approaches’ ability to capture a comprehensive view of the network, reducing their effectiveness against sophisticated attacks.
@@
Multimodal Federated Learning. Multimodal FL integrates heterogeneous data sources, such as images, audio, and time-series signals, to capture complementary information and improve collaborative learning robustness. While this paradigm is gaining traction in fields such as healthcare, human activity recognition, and autonomous systems, its application to wireless security remains limited [mmsurvey]. Despite its potential, existing multimodal FL approaches across domains still face two significant limitations. Firstly, they exhibit poor performance under non-IID conditions, an inherent challenge in real-world deployments where clients experience diverse environments and modality distributions [chen2022towards]. Secondly, they frequently overlook the increased computational cost of processing multiple data streams, often assuming that all clients have uniform and sufficient resources [Che2023Multimodal]. These limitations hinder their applicability in resource-constrained, heterogeneous edge scenarios, such as wireless jamming detection.
@@ III-C Training Workflow
FedJam adopts a synchronous FL setup to collaboratively train a multimodal model across multiple edge clients. Each client performs local updates on its private spectrogram and KPI data, and periodically transmits selected model parameters, namely those of the spectrum encoder, KPI encoder, and classifier, to a central server. To minimize communication costs, these updates may be compressed before transmission. The server aggregates the received updates using any aggregation method (e.g., FedAvg [fedavg]) and broadcasts the updated global model back to all clients. This process is repeated for a fixed number of communication rounds or until convergence, enabling the multimodal model to learn effectively from decentralized data without sharing raw samples.
@@ IV-A Experimental Environment
Network Connectivity. Connectivity between the edge devices and the central server is supported through both Wi-Fi and 5G. Wi-Fi connectivity is provided by a TP-Link Archer C7 router running OpenWRT firmware, operating exclusively in the 2.4 GHz band across the three non-overlapping 20 MHz channels (1, 6, and 11). 5G connectivity is enabled through a fully customized private 5G network operating in the n78 TDD band with 20 MHz bandwidth centered at 3.319 GHz, using a 5 ms radio frame structure with 7 downlink slots and 2 uplink slots at 30 kHz subcarrier spacing. This private 5G deployment is supported by a separate testbed developed in-house, shown in Fig. 3(b) and Fig. 3(c), featuring a monolithic base station (gNB), connected to a USRP N300 SDR via a 10 Gbps fronthaul link, with the SDR further connected to 6 dBi directional antennas to enable over-the-air transmission. The gNB is further connected to the 5G core network via a dedicated 10 Gbps backhaul link, with all 3GPP-compliant Virtual Network Functions (VNFs) deployed as containerized services within a customized Kubernetes cluster. Both the Radio Access Network (RAN) and core network VNFs are implemented using the latest release of the OpenAirInterface (OAI) 5G software stack [oai].
@@ V-A Baselines
  • FewShotDense: A unimodal spectrogram-based convolutional neural network built on DenseNet121, previously used for few-shot interference classification tasks [federated_learning_spectrogram_5].
  • SpectroNet: A unimodal EfficientNet-B0-based architecture trained on spectrogram data for multi-class jamming detection and classification [spectrogram_1].
  • JamShield: A unimodal LSTM-based architecture trained on cross-layer KPI time series data for multi-class jamming detection [jamshield].
@@
  • Notable Unimodal Models: We also experiment with two lightweight unimodal architectures to assess individual modality performance under constrained resources: (i) an 1D-CNN [kiranyaz20211d] operating on KPI time series and (ii) TinyViT [wu2022tinyvit], a compact vision transformer applied to spectrogram inputs.
@@ V-B FedJam Configuration
For spectral feature extraction, the Spectrum Encoder employs MobileNetV3-Small [howard2019searching], a lightweight convolutional neural network pretrained on ImageNet, producing a 1024-dimensional feature vector that captures localized interference patterns in the frequency domain. The KPI Encoder is implemented using a compact Transformer-based architecture composed of four self-attention layers, preceded by a linear projection to 128 dimensions. The resulting sequence is aggregated using an attention-based pooling mechanism to produce a 128-dimensional embedding summarizing the network’s temporal performance. The Fusion Module combines the two modality-specific embeddings via concatenation, forming a 1152-dimensional joint representation. This fused vector is then passed to the Multimodal Projection Head Module, implemented as a two-layer fully connected neural network with 256 hidden units and ReLU activation, followed by an output layer producing logits over the four jamming classes. Finally, all of FedJam’s components are jointly fine-tuned end-to-end using our collected multimodal dataset.
@@ V-D Evaluation Results
For our simulation experiments, we utilized the Flower framework [beutel2020flower], a widely adopted and flexible platform in the FL community that enables evaluation of federated solutions across diverse scales and heterogeneous settings.
@@
FewShotDense [federated_learning_spectrogram_5] 95.10 $\pm$ 0.31 88.10 $\pm$ 0.24 84.10 $\pm$ 0.39
SpectroNet [spectrogram_1] 96.00 $\pm$ 0.27 93.80 $\pm$ 0.34 89.00 $\pm$ 0.25
JamShield [jamshield] 58.40 $\pm$ 0.40 50.70 $\pm$ 0.21 46.70 $\pm$ 0.28
1D-CNN [kiranyaz20211d] 46.60 $\pm$ 0.33 42.53 $\pm$ 0.30 40.89 $\pm$ 0.29
TinyViT [wu2022tinyvit] 89.60 $\pm$ 0.87 84.60 $\pm$ 0.52 77.10 $\pm$ 0.92
@@ 10 Clients 25 Clients 50 Clients
In addition to fusion strategies, the choice of aggregation algorithm can influence convergence and robustness in FL. We evaluate three widely used aggregation methods: FedAvg, FedOpt [fedopt], and FedProx [fedprox], to assess their compatibility with our framework. As shown in Fig. 7(b), all algorithms perform similarly, with FedAvg offering a marginal improvement. These results highlight that our method does not rely on a specific aggregation scheme. Indeed, FedJam is designed to be modular and aggregation-agnostic, enabling seamless integration with future or domain-specific optimizers, and supporting flexible experimentation in practical deployments.
@@
We additionally evaluate the total training time required under different device energy modes, where all participating clients operate at the same power level. The Jetson Orin Nano devices support three energy modes: Mode 0 (15W), Mode 1 (25W), and Mode 2 (MAXN SUPER), which progressively scale CPU/GPU frequencies and memory bandwidth [energy_modes]. As shown in Fig. 11, higher energy modes consistently reduce training time across both Wi-Fi and 5G. For instance, under Wi-Fi with 4 clients, total training time decreases from 575 minutes (15W) to 451 minutes (MAXN SUPER), corresponding to a 21.6% reduction. Similarly, with 16 clients, training time drops from 290 to 220 minutes—a 24.1% improvement. These improvements reflect lower per-device training time due to increased hardware utilization and faster computation, although diminishing gains are observed beyond 8 clients due to reduced per-client workload.